OpenStudy (anonymous):

A group of ten women and twelve men must select a three person committee. How many committees of one woman and two men are possible? --- This is a counting method and probability problem.

OpenStudy (anonymous):

10 * 12C2 = 10 * (12*11/2)= 10 * 66 = 660

OpenStudy (anonymous):

seems a heck of a lot !!

OpenStudy (anonymous):

but i think its correct as there any of ten women combuned with the possible combinations of two from 12 men

OpenStudy (anonymous):

why you divided 11 by 2 ?

OpenStudy (anonymous):

Think about it

OpenStudy (anonymous):

i multiplied 12 by 11 to give 32 then divided by = 132 / 2 = 66

OpenStudy (anonymous):

i think the cmbinations of two man are just 132 = 12*11

OpenStudy (anonymous):

ok ok forget it

OpenStudy (anonymous):

Yeah I think @AlleluiaSB is correct on this one. Thank you both.

OpenStudy (anonymous):

the formula for combinations of r from n items = n! / (n-r)! r! which is the same thing as for example 2 combinations in 4 = 4*3 / 2*1

OpenStudy (anonymous):

12*11 = number of permutations not combinations - in permutations the order is important in combinations it is not

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

in my calculations some groups had the same couple of man. sorry about it.

OpenStudy (anonymous):

for example take a b and c permutations of 2 are ab ac bc ba ca cb combinations are ab ac bc - the order does not matter

OpenStudy (anonymous):

yep - no probs

OpenStudy (anonymous):

what does 10 * 12C2 mean?